OpenStudy (anonymous):

Which of the following is a possible set of quantum numbers for an electron? (1, 1, 0, +½) (2, 1, 2, +½) (3, 2, 0, -½) (3, -2, 1, -½)

thomaster (thomaster):

Do you know What kind of quantum numbers those 4 are?

OpenStudy (anonymous):

._. What do you mean?

thomaster (thomaster):

I mean do you know what the first, second, third and fourth number mean?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I Kind Of Forgot Like the rule of it though.

thomaster (thomaster):

What's the first number?

OpenStudy (anonymous):

-_- N, The Principal Quantum Number.

OpenStudy (anonymous):

L, the azimuthal number, Ml the magnetic, and the last is the spin. I Just need to know the part where it says which numbers can be in each levels...if that makes sense.

thomaster (thomaster):

N ranges from 1 to infinity L ranges from 0 to N-1 Ml ranges from -L to L Ms is either \(-\frac{1}2\) or \(\frac{1}2\)

thomaster (thomaster):

to check which one is correct, you have to look at N, then look at the possibilities for L For example in the first one: N = 1 L ranges from 0 to N-1, N-1 = 1-1 = 0 So L can only be 0. In the first option, L=1 so the first one is not a possible set of quantum numbers

OpenStudy (anonymous):

I think it's the second one?

thomaster (thomaster):

Second one: N=2 L= 0 \(\to\) N-1 L=2-1=1 So L ranges from 0 to 1 That means L for the second option is possible. Now the magnetic number Ml = -L to +L = -1 to +1 It's 2 in the option so second one is also not possible

thomaster (thomaster):

Now check if the third one is possible, can you do that? :)

OpenStudy (anonymous):

Ohhh Yea I Got It Thank You!

OpenStudy (anonymous):

Wait, What about the spin?

thomaster (thomaster):

The spin is always correct. It can only be -1/2 or +1/2